# Nemesis: A Multi-Scale, Multi-Physics Algorithm for Astrophysics

### 03 Mar 2026 v1

@@ 5 Conclusions
After describing the motivation and theory behind the scheme in section 2, a suite of simulations are run to test its performance. The first test considers asteroids within a star cluster evolution. When comparing results with a direct $N$-body code, the final orbital parameters of asteroids between algorithms were found to be statistically indistinguishable. When tracking the energy error over longer periods of time, Nemesis successfully suppresses energy drift when using symplectic codes. This feature is essential for simulating systems over secular timescales. The success in modelling secular effects is highlighted by the second set of tests with its ability to reproduce the ZLK oscillations in hierarchical triples.

### 10 Mar 2026 v2

@@ 5 Conclusions
After describing the motivation and theory behind the scheme in section 2, a suite of simulations are run to test its performance. The first test considers asteroids within a star cluster’s evolution. When comparing results with a direct $N$-body code, the final orbital parameters of asteroids between algorithms were found to be statistically indistinguishable. When tracking the energy error over longer periods of time, Nemesis successfully suppresses energy drift when using symplectic codes. This feature is essential for simulating systems over secular timescales. The success in modelling secular effects is highlighted by the second set of tests, with its ability to reproduce the ZLK oscillations in hierarchical triples.